# Unified Domain Adaptive Semantic Segmentation

### v2

@@ IV-D Further Analysis
| , = 1.00, 0.10 | 55.20 |
@@ IV-D1 Hyperparameter Sensitivity
For model optimization, we utilize two hyperparameters, in Eq. (22) and Eq. (24) and in Eq. (23). In Table IX, we progressively increase the values of and using Viper Cityscapes-Seq as an example to explore the hyperparameter sensitivity. The model achieves an mIOU of 56.16 when these parameters are set to 0.20 and 0.01 and a suboptimal mIOU of 55.20 when the two parameters are set to 1.00 and 0.10. The results display slight fluctuations with parameter setting changes, showing our method’s robustness. For convenience, we set and to 1 for another Synthea-Seq Cityscapes-Seq benchmark.
@@ IV-D2 Results with Model Initialization
As shown in Table X, we achieve 59.62 and 54.94 mIOU on two benchmarks by end-to-end training, surpassing previous methods trained both end-to-end and in stages. Initializing our models with warm-up parameters from previous works in the “Original” column can further enhance adaptation. Notably, CMOM [39] initializes model parameters and generates initial pseudo-labels offline using a warm-up model trained in [6]. In contrast, we solely undergo model initialization. The above analysis underscores the remarkable adaptation of our method on UDA-VSS.
| Modules | SYN-SeqCity-Seq | VIPERCity-Seq | ||||||
| Src. | SSL | QuadMix | Agg. | Time (s/iter) | GPU memory | Time (s/iter) | GPU memory | |
| ✓ | 1.40 | 8.78 G | 1.38 | 8.78 G | ||||
| ✓ | 2.31 | 12.52 G | 2.34 | 13.54 G | ||||
| ✓ | ✓ | 3.78 | 21.24 G | 3.99 | 21.90 G | |||
| ✓ | ✓ | ✓ | 4.26 | 24.92 G | 4.37 | 25.90 G | ||
| ✓ | ✓ | ✓ | 5.64 | 29.11 G | 5.89 | 29.85 G | ||
@@ IV-D3 The Throughput Measured for Different Modules
In Table XI, we present the training time and GPU memory footprint associated with the integration of different components. “Src.” refers to training only with the source domain data. “” corresponds to the ablation setting ID 11 in Table VI. Compared to self-supervised learning “SSL”, the integration with QuadMix and leads to a modest increase in training time and GPU memory, and “Agg” exhibits relatively lighter overhead.
@@ IV-D5 Results of Different Training Iterations
In Fig. 10, we provide a more comprehensive presentation of semantic segmentation results of our models trained at different iterations. Notably, the best models for both two benchmarks emerge between the 35k and 40k iterations, providing a clearer insight into the evolution and stability of the training process.
@@ V Conclusion
This paper tackles the more challenging task of unified UDA-SS, addressing both video and image scenarios in a unified manner. Further, we propose QuadMix to address fragmentary and insufficient feature gap bridging caused by distinct point attributes. Specifically, we propose quad-directional intra- and inter-domain mixing at both pixel and feature levels, aiming to generalize intra-domain, enhance inter-domain, and alleviate feature inconsistencies. Additionally, we propose flow-guided video feature aggregation for fine-grained domain alignment, applicable to image scenarios as well. Extensive experiments on four benchmarks demonstrate our method’s superiority across video and image scenarios. One limitation is that while quad-mixed samples enhance model adaptation, some unrealistic samples may mislead the model. Future work will explore generation-based reliable mixing for open-vocabulary domain adaptation using vision-language models.
@@ Acknowledgment
This work is supported in part by the Major Program of the National Natural Science Foundation of China (NSFC) under Grants 61991400, and 61991404, in part by the NSFC under Grants 62103092, and Program of China Scholarship Council 202306080142. The authors appreciate Chunhua Shen and Lan Deng’s helpful suggestions.

### v3

@@ IV-C3 Ablation Study on Flow-Guided Spatio-Temporal Feature Aggregation Module
| , = 1.00, 0.10 | 55.21 |
| Modules | SYN-SeqCity-Seq | VIPERCity-Seq | |||||
| Src. | QuadMix | QuadMix* | Agg. | Time (s/iter) | GPU memory | Time (s/iter) | GPU memory |
| ✓ | 1.40 | 8.78 G | 1.38 | 8.78 G | |||
| 2.31 | 12.52 G | 2.34 | 13.54 G | ||||
| ✓ | 3.78 | 21.24 G | 3.99 | 21.90 G | |||
| ✓ | ✓ | 4.26 | 24.92 G | 4.37 | 25.90 G | ||
| ✓ | ✓ | 5.64 | 29.11 G | 5.89 | 29.85 G | ||
IV-D Further Analysis
We conduct further analysis on the general task of video UDA-SS, using the DeepLab-V2 backbone by default.
@@ IV-D1 Hyperparameter Sensitivity
For model optimization, we utilize two hyperparameters, in Eq. (22) and Eq. (24) and in Eq. (23). In Table IX, we progressively increase the values of and using Viper Cityscapes-Seq as an example to explore the hyperparameter sensitivity. The model achieves an mIOU of 56.16 when these parameters are set to 0.20 and 0.01 and a suboptimal mIOU of 55.21 when the two parameters are set to 1.00 and 0.10. The results display slight fluctuations with parameter setting changes, showing our method’s robustness. For convenience, we set and to 1 for another Synthea-Seq Cityscapes-Seq benchmark.
@@ IV-D2 Results with Model Initialization
As shown in Table X, we achieve 59.62 and 54.94 mIOU on two benchmarks by end-to-end training, surpassing previous methods trained both end-to-end and stage-wise. Initializing our models with warm-up parameters from previous works in the “Original” column can further enhance adaptation. Notably, CMOM [2] initializes model parameters and generates initial pseudo-labels offline using a warm-up model trained in [16]. In contrast, we solely undergo model initialization. The above analysis underscores the remarkable adaptation of our method on video UDA-SS.
@@ IV-D3 The Throughput Measured for Different Modules
In Table XI, we present the training time and GPU memory footprint associated with the integration of different components. “Src.” refers to training only with the source domain data. “QuadMix*” corresponds to the ablation setting ID 11 in Table VI. Compared to “Src.”, the proposed QuadMix and leads to a moderate increase in training time and GPU memory, and “Agg” exhibits relatively lighter overhead.
@@ IV-D5 Results of Different Training Iterations
In Fig. 11, we provide a more comprehensive presentation of the semantic segmentation results of our models trained at different iterations. Notably, the best models for both two benchmarks emerge between the 35k and 40k iterations, providing a clearer insight into the evolution and stability of the training process.
@@ V Conclusion
This paper tackles the more challenging task of unified UDA-SS across both video and image scenarios by presenting an efficient and comprehensive approach. Specifically, we propose the QuadMix that involves quad-directional intra- and inter-domain mixing at both pixel and feature levels, aiming to generalize intra-domain features, enhance inter-domain gap bridging, and alleviate feature inconsistencies for domain invariant learning. Furthermore, we propose flow-guided video feature aggregation for fine-grained domain alignment, applicable to image scenarios as well. Our method is compatible with both CNN and transformer-based architectures. Extensive experiments on four challenging benchmarks demonstrate the superiority of our method.
Limitation: One limitation is that, while the quad-mixing enhances domain adaptation effectively, the theoretical interpretability of underlying mechanisms remains to be explored.
Future work: An interesting aspect of future work is to explore unified techniques across video and image scenarios, particularly in multi-modal domain adaptation, open-set domain adaptation, and extensions to instance or panoramic segmentation, as well as pixel-level anomaly detection.
@@ Acknowledgment
This work is supported in part by the Major Program of the National Natural Science Foundation of China (NSFC) under Grants 61991404, in part by the NSFC under Grants 62103092, and Program of China Scholarship Council 202306080142. The authors appreciate the valuable suggestions of Chunhua Shen and Lan Deng.